OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. 0 = x2 + 5x + 6 0 = x2 + 4x + 4 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 4x + 4 different from y = x2 + 5x + 6?

OpenStudy (anonymous):

x2+5x+6 = (x+3)(x+2) = -3,-2 x2+4x+4 = (x+2)^2 = -2 They both have a solution of -2 and 2 real solutions. They both have a different vertex and different positions on the coordinate plane.